OpenStudy (anonymous):

Solve 3x – 5y = 15 for y. y = -5/3x – 3 y = 5/3x + 15 y = -3/5x + 15 y = 3/5x – 3

OpenStudy (nurali):

3x – 5y = 15 3x-15=5y y=3x-15/5 y=3x/5-15/5 y=3x/5-3

OpenStudy (anonymous):

thanks for actually explaining

OpenStudy (nurali):

My Pleasure.